# Broadly Tunable Quantum Enhanced Raman Microscopy for Advancing Bioimaging

### v1

@@
Broadly Tunable Quantum Enhanced Raman Microscopy for Advancing Bioimaging
††journal: opticajournal††articletype: Research ArticleStimulated Raman scattering (SRS) microscopy has emerged as a powerful technique for probing the spatiotemporal dynamics of molecular bonds with exceptional sensitivity, resolution, and speed. However, classically, its performance remains fundamentally constrained by optical shot noise, which imposes a strict limit on detection sensitivity and speed. Here, we demonstrate a quantum-enhanced SRS microscopy platform that circumvents this barrier by harnessing amplitude-squeezed light. Specifically, we generate a Stokes beam with of amplitude squeezing using traveling-wave optical parametric amplification in second-order nonlinear waveguides, and combine it with a tunable coherent pump to access vibrational modes spanning from to . Applied to quantum imaging of metabolites in biological tissue (pork muscle), our quantum-enhanced Raman microscope achieves an average noise suppression of and a enhancement in signal-to-noise ratio (SNR) – to the best of our knowledge, the largest improvement reported to date in quantum-enhanced SRS microscopy of biological samples.
@@ 1 Introduction
Raman-based microscopy provides a powerful approach for probing molecular vibrations with high chemical specificity in a label-free and non-destructive manner [10]. However, the intrinsically weak nature of spontaneous Raman scattering limits both its sensitivity and acquisition speed, particularly in live-cell and clinical imaging contexts. To overcome these challenges, stimulated Raman scattering (SRS) microscopy was developed as a technique that resonantly drives molecular vibrations using synchronized pump and Stokes beams whose frequency difference is tuned to a specific vibrational mode of interest. By amplifying Raman signals by several orders of magnitude while preserving a linear dependence on molecular concentration, SRS enables quantitative, high-speed, and background-free chemical imaging [11, 12]. These capabilities have established SRS as a robust technique for spatially and chemically resolved imaging of lipids, proteins, metabolites, and pharmaceutical compounds in living systems, as well as for real-time biomedical applications, including intraoperative diagnostics [13, 14].
Extending the principles of quantum-enhanced sensing to Raman-based microscopy provides a direct means to overcome the shot-noise limit that currently constrains SRS sensitivity. In conventional implementations, detection can be improved by increasing optical power or extending integration time, but for fragile biological tissues, excitation must remain below photodamage thresholds to preserve their native dynamics [15, 16]. By employing squeezed probe states, which intrinsically exhibit noise levels below the standard shot-noise limit, quantum-enhanced Raman microscopy achieves higher sensitivity and faster acquisition, hereby reduced photodamage [17].
@@
One of the first experimental proof-of-concept demonstrations of quantum-enhanced SRS (QE-SRS) employed continuous-wave amplitude-squeezed light, achieving a noise reduction of 3.6 dB at 2950 cm-1 in polymer samples [18]. Subsequently, pulsed amplitude squeezing enabled sub-shot-noise imaging of molecular bonds within live yeast cells, yielding a 1.3 dB noise reduction at 2850 cm-1 [19]. Additionally, a quantum-enhanced balanced detection (QE-BD) scheme extended operation into the high-power regime (>30 mW), demonstrating a 3.11 dB noise reduction compared to classical detection strategies [20]. Beyond sensitivity improvements, squeezed light has also been leveraged to enhance imaging capabilities [21], while integration of bright squeezed probe beams into live-cell SRS microscopy enabled 20% faster imaging with reduced photodamage [22]. All these studies demonstrate remarkable improvements offered by QE-SRS compared to conventional implementations. However, their investigation range has so far been restricted to a narrow bandwidth of Raman vibrational modes, primarily within the functional group region (2800–3100 cm-1). This limitation restricts the extraction of molecular information, as the Raman signatures of lipids, proteins, and nucleic acids exhibit substantial spectral overlap within this vibrational region. By extending access into the fingerprint region (500–1800 cm-1), it becomes possible to resolve these overlapping contributions and unambiguously distinguish between the vibrational signatures of different biomolecular species, thereby providing a far richer and more specific chemical characterization of biological tissues.
@@
In this work, we present a quantum-enhanced SRS microscope that employs a bright amplitude-squeezed Stokes beam as the probe field. The squeezed beam provides more than 5 dB of noise reduction while maintaining an optical power of 3.75 mW, making it compatible with picosecond SRS imaging. Integrated into a tunable SRS platform, it enables broadband operation across both the fingerprint region (1450–1650 cm-1) and the CH-stretch region (2800–3100 cm-1), allowing chemically specific mapping of polystyrene reference samples and biological tissue with enhanced sensitivity. Whereas previous quantum-enhanced SRS demonstrations were confined to the CH-stretch region, the present work achieves, for the first time, quantum-enhanced vibrational imaging spanning both spectral regions, thereby enabling chemically distinguishable imaging in biologically relevant environments.
@@ 2.1 Full scheme
Following interaction with the sample, the Raman pump and Stokes beams are collected by the second MO 50. The Stokes beam is then separated from the Raman pump using a combination of a dichroic mirror (DM, Thorlabs DMSP1020B) and a long-pass filter (LPF, Thorlabs FELH1000) and directed onto a custom-built photodetector (InGaAs PD), adapted from the design of Ref. [19]. The detector uses an InGaAs PIN Hamamatsu G10899-003K photodiode with a measured quantum efficiency of 97 %. The detector is resonantly tuned to amplify signals at the 19.3 MHz modulation frequency. The SRS signal is acquired using a digital lock-in amplifier implemented on the Moku:Lab platform (MOKU, Liquid Instruments), with a 12-bit vertical resolution, and detection bandwidth of 2 kHz around the demodulation frequency achieving low-noise signal recovery. In addition, the Moku:Lab is used to generate the 19.3 MHz drive signal for the EOM, which enables precise amplitude modulation of the pump beam for SRS detection. For synchronization between the XYZ scanning stage and data acquisition, we implement a custom-made Python script, achieving a scan rate of 15 data points per second. While sufficient for proof-of-concept measurements, this rate is modest compared to commercial high-speed SRS systems. The limitation arises from the inherent speed constraints of our mechanical scanning stage, which are slower than resonant scanners or galvo mirrors, and from potential bottlenecks in data transfer and storage associated with our Python-based acquisition.
@@ 2.3 Tunability of QE-SRS microscope
We investigated three of the most prominent Raman lines of polystyrene: 3050 cm-1, corresponding to the aromatic C–H stretching mode; 1600 cm-1, associated with the skeletal C=C stretching mode of the phenyl ring; and 1003 cm-1, corresponding to the symmetric ring-breathing mode of the phenyl group. To match the targeted Raman lines, the pump wavelength from the picoEmerald laser was tuned from 803.5 nm to 961 nm, covering the spectral range from the C–H stretching mode to the ring-breathing mode. The Raman pump power was set to approximately 34 mW, while the coherent Stokes beam power was maintained at around 3.75 mW to ensure a sufficient signal level. At each scanning point, the signal was averaged over 1024 acquisitions, resulting in a dwell time of around 66 ms per pixel. The SRS images obtained with our home-built microscopy system show good agreement with the corresponding white-light wide-field image, confirming both the spatial resolution and spectral tunability of the system.
@@ 3.1 QE-SRS on a pork muscle tissue
The Raman band at 2940 cm-1 originates from CH2/CH3 stretching modes present in both proteins and lipids and therefore serves as an indicator of overall protein–lipid content in tissue. The measurement was performed using an average optical power of 32 mW for the Raman pump and 3.75 mW for the Stokes beam focused on the sample. The white-light wide-field image is shown in Fig. 3(a). Spatial mapping of the muscle structure was achieved via raster scanning with a step size of 750 nm, with signal averaging over 1024 acquisitions per pixel. The measured SRS signal (Fig. 3(b)) clearly reproduces the structural features of the muscle tissue. Fig. 3(d),(e) further illustrate the enhancement in SNR achieved by employing amplitude-squeezed light. In the case of SRS gain detection, the SNR is [18]
where is the SRS signal intensity, and are pump and Stokes beam intensities, respectively, and is the variance of the detected Stokes beam intensity that quantifies the noise level. By using amplitude-squeezed light, we are able to surpass the shot-noise level, thereby enhancing the SNR in proportion to the inverse square root of the noise reduction factor. In Fig. 3(d) and (e), we observed an SNR improvement of approximately 43% across the entire investigated area (m2), correspond to a noise reduction of compared to the coherent Stokes beam configuration. The initial 5.2 dB of squeezing was reduced to 3.3 dB at the detector due to 24% optical loss arising from optics and sample transmission.
In addition, we investigated the scaling behavior of the SNR with respect to the Raman pump power. As discussed above, the SNR is expected to scale linearly with pump power. To validate this, we measured the SNR of the SRS signal within the white square region shown in Fig. 3(b). The results are presented in Fig. 3(c). Here, the SNR was averaged over a m2 area, while the pump power was varied from 12 mW to 49 mW with the Stokes beam power fixed at 3.75 mW. The blue and red dashed lines correspond to linear fits of the acquired data for coherent (red trace) and amplitude-squeezed (blue trace) Stokes beams, respectively. The increased slope observed in the squeezed-light case corresponds to an estimated noise reduction of approximately 3.3 dB, yielding a 43% improvement in SNR, which is consistent with the enhancements observed in panels (d) and (e).
@@ 3.2 Multifrequency biological imaging
To demonstrate chemical specificity, we performed SRS imaging at four vibrational bands characteristic of proteins and lipids. Explicitly, we investigated several prominent Raman vibrational modes: (general C–H stretching in proteins, lipids, and nucleic acids), widely used as a marker of total protein and lipid content; (symmetric CH2 stretching in lipids); (amide I band, primarily C=O stretching in proteins), a key indicator of protein content and secondary structure; and (CH2/CH3 bending in lipids and proteins), often used as a marker for lipid–protein balance. Together, these four bands give a complementary biochemical fingerprint of muscle tissue. For these measurements, the Raman pump was tuned to , , , and with corresponding powers at the sample of 32 mW, 32 mW, 30 mW, and 38 mW, respectively. The Stokes beam power was maintained at in all cases. However, systematically higher SNR were observed at the CH-stretch bands (2850 and 2940 cm-1) compared to the fingerprint modes. This enhancement is attributed to the larger Raman cross-sections and the higher abundance of CH groups in tissue lipids and proteins, particularly in lipid-rich structures such as cell membranes and adipose tissue. In contrast, Fig. 4 shows regions with SRS signals at 1450 cm-1 and 1650 cm-1, where the CH-stretch bands at 2850 and 2940 cm-1 are absent. This spectral pattern indicates regions with elevated protein content and minimal lipid contribution, consistent with lean myofibrillar or connective tissue in muscle [26, 27]. The results presented in Fig. 4 demonstrate a consistent enhancement in SNR across all examined Raman shifts when utilizing the amplitude-squeezed Stokes beam. A noise reduction of dB was achieved across all measured vibrational modes, corresponding to a improvement in SNR compared to the classical (coherent) Stokes configuration. Fig. 4(f) and (g) show cross-sections of the SNR maps. These cross-sections clearly illustrate the improvements achieved with squeezed light compared to coherent illumination. This consistent enhancement underscores the robustness of quantum-assisted amplification in improving Raman signal quality. The observed gains highlight the potential of QE-SRS to extend the sensitivity of vibrational imaging in biological tissues, enabling more precise and reliable real-time biochemical analysis, even at lower optical powers or in samples with weak Raman signatures.
@@ 4 Conclusion
We validated the performance of the quantum-enhanced SRS microscope by conducting measurements on both benchmark polystyrene samples and biological samples (pork muscle tissue). The biological measurements revealed an average noise reduction of 3.6 dB, corresponding to a 51% improvement in SNR compared to classical detection. To the best of our knowledge, this represents the highest reported enhancement in QE-SRS microscopy applied to biological samples. Importantly, the results confirm that quantum squeezing can be practically integrated into a picosecond-pulsed SRS platform and translated from proof-of-concept demonstrations on simple materials to more complex biological tissue environments. Along with strong squeezing, we also report, for the first time, QE-SRS across both regions in the fingerprint and high-wavenumber CH-stretch regions. Our approach provides quantum SNR enhancement in both regions, enabling sensitive and label-free imaging and identification of important biomarkers, such as proteins, lipids, amid and nucleic acids in biological samples.
Looking forward, further improvements in coupling efficiency, waveguide engineering, and pump power stabilization are expected to increase the achievable squeezing levels and long-term stability of the source. Integration with more advanced lock-in detection schemes and fast-scanning microscopes could enable real-time quantum-enhanced imaging of living biological samples, extending the method from proof-of-concept demonstrations towards practical biomedical applications. Moreover, combining squeezed-light excitation with computational imaging and machine learning approaches may open new possibilities for super-resolved Raman microscopy with reduced photo damage and enhanced molecular contrast.

### v2

# Broadly tunable quantum-enhanced Raman microscopy for advancing bioimaging

@@
Abstract
Stimulated Raman scattering (SRS) microscopy has emerged as a powerful technique for probing the spatiotemporal dynamics of molecular bonds with exceptional sensitivity, resolution, and speed. However, classically, its performance remains fundamentally constrained by optical shot noise, which imposes a strict limit on detection sensitivity and speed. Here, we demonstrate a quantum-enhanced SRS microscopy platform that circumvents this barrier by harnessing amplitude-squeezed light. Specifically, we generate a Stokes beam with of amplitude squeezing using traveling-wave optical parametric amplification in second-order nonlinear waveguides, and combine it with a tunable coherent pump to access vibrational modes spanning from to . Applied to quantum imaging of metabolites in biological tissue (pork muscle), our quantum-enhanced Raman microscope achieves an average noise suppression of and a enhancement in signal-to-noise ratio (SNR) – to the best of our knowledge, the largest improvement reported to date in quantum-enhanced SRS microscopy of biological samples.
@@ 1 Introduction
Raman-based microscopy provides a powerful approach for probing molecular vibrations with high chemical specificity in a label-free and non-destructive manner [6]. However, the intrinsically weak nature of spontaneous Raman scattering limits both its sensitivity and acquisition speed, particularly in live-cell and clinical imaging contexts. To overcome these challenges, stimulated Raman scattering (SRS) microscopy was developed as a technique that resonantly drives molecular vibrations using synchronized pump and Stokes beams whose frequency difference is tuned to a specific vibrational mode of interest. By amplifying Raman signals by several orders of magnitude while preserving a linear dependence on molecular concentration, SRS enables quantitative, high-speed, and background-free chemical imaging [18, 9]. These capabilities have established SRS as a robust technique for spatially and chemically resolved imaging of lipids, proteins, metabolites, and pharmaceutical compounds in living systems, as well as for real-time biomedical applications, including intraoperative diagnostics [13, 23]. Extending the principles of quantum-enhanced sensing to Raman-based microscopy provides a direct means to overcome the shot-noise limit that currently constrains SRS sensitivity. In conventional implementations, detection can be improved by increasing optical power or extending integration time, but for fragile biological tissues, excitation must remain below photodamage thresholds to preserve their native dynamics [26]. By employing squeezed probe states with noise levels below the standard shot-noise limit, quantum-enhanced Raman microscopy achieves enhanced sensitivity and faster acquisition rates, thereby minimizing photodamage in biological samples [22]. Beyond Raman-based approaches, close to 4 dB of quantum enhancement has been demonstrated in continuous-wave (CW) Brillouin microscopy, derived from small regions within the scanned area, where two-mode intensity-difference squeezed light generated via four-wave mixing in a Rb vapor cell was used to probe the mechanical properties of biological samples [16]. While Raman spectroscopy provides chemically specific biochemical contrast and is generally more suitable for bioimaging, Brillouin microscopy delivers complementary information on the viscoelastic and mechanical properties of the sample.
@@
One of the first experimental proof-of-concept demonstrations of quantum-enhanced SRS (QE-SRS) employed CW amplitude-squeezed light, achieving a noise reduction of 3.6 dB at 2950 cm-1 in polymer samples [7]. Subsequently, pulsed amplitude squeezing enabled sub-shot-noise imaging of molecular bonds within live yeast cells, yielding a 1.3 dB noise reduction at 2850 cm-1 [5]. Additionally, a quantum-enhanced balanced detection (QE-BD) scheme extended operation into the high-power regime (>30 mW), demonstrating a 3.11 dB noise reduction compared to classical detection strategies [25]. Beyond sensitivity improvements, squeezed light has also been leveraged to enhance imaging capabilities [11], while integration of bright squeezed probe beams into live-cell SRS microscopy enabled 20% faster imaging with reduced photodamage [22]. All of these studies demonstrate remarkable improvements offered by QE-SRS compared to conventional implementations. However, previous studies have been limited to a narrow subset of Raman vibrational modes. In Refs. [22, 5], measurements were confined to the CH-stretch region, where overlapping Raman peaks from multiple biomolecules (e.g., lipids and fats) reduce chemical specificity. Similarly, Ref. [7] investigated simple polymer samples using a continuous-wave (CW) SRS configuration, which further restricts the extraction of detailed molecular information. Within the CH-stretch region, the Raman signatures of lipids, proteins, and nucleic acids exhibit substantial spectral overlap, making it difficult to discriminate between them. By extending measurements into the fingerprint region (500–1800 cm-1), these overlapping contributions can be resolved, allowing unambiguous identification of distinct biomolecular vibrations and enabling a far richer and more specific chemical characterization of biological tissues.
@@
In this work, we present a quantum-enhanced SRS microscope that employs a bright amplitude-squeezed Stokes beam as the probe field. The squeezed beam provides more than 5 dB of noise reduction while maintaining an optical power of 3.75 mW, making it compatible with picosecond SRS imaging. Integrated into a tunable SRS platform, it enables broadband operation across both the fingerprint region (1450–1650 cm-1) and the CH-stretch region (2800–3100 cm-1), allowing chemically specific mapping of polystyrene reference samples and biological tissue with enhanced, achieving an approximately three-orders-of-magnitude improvement in material sensitivity compared to the CW approach reported in Ref. [7]. Whereas previous quantum-enhanced SRS demonstrations were confined to the CH-stretch region, the present work achieves, for the first time, quantum-enhanced vibrational imaging spanning both spectral regions, thereby enabling chemically distinguishable imaging in biologically relevant environments.
@@ 2.1 Full scheme
Following interaction with the sample, the Raman pump and Stokes beams are collected by the second MO 50. The Stokes beam is then separated from the Raman pump using a combination of a dichroic mirror (DM, Thorlabs DMSP1020B) and a long-pass filter (LPF, Thorlabs FELH1000) and directed onto a custom-built photodetector (InGaAs PD), adapted from the design of Ref. [5]. The detector uses an InGaAs PIN Hamamatsu G10899-003K photodiode with a measured quantum efficiency of 97 %. The detector is resonantly tuned to amplify signals at the 19.3 MHz modulation frequency. The SRS signal is acquired using a digital lock-in amplifier implemented on the Moku:Lab platform (MOKU, Liquid Instruments), with a 12-bit vertical resolution, and detection bandwidth of 2 kHz around the demodulation frequency achieving low-noise signal recovery. In addition, the Moku:Lab is used to generate the 19.3 MHz drive signal for the EOM, which enables precise amplitude modulation of the pump beam for SRS detection. For synchronization between the XYZ scanning stage and the data acquisition, we implement a custom Python-based control script, achieving a scan rate of 15 data points per second. While sufficient for proof-of-concept measurements, this rate is modest compared to commercial high-speed SRS systems. The Moku acquisition window, consisting of 1024 points, corresponds to approximately 10 ms. Therefore, the limitation arises primarily from the need to synchronize data acquisition with stage positioning, with the dwell time per pixel (66 ms) being largely dominated by communication and data handling overhead in the Python-based acquisition framework.
@@ 2.3 Tunability of QE-SRS microscope
We investigated three of the most prominent Raman lines of polystyrene: 3050 cm-1, corresponding to the aromatic C–H stretching mode; 1600 cm-1, associated with the skeletal C=C stretching mode of the phenyl ring; and 1003 cm-1, corresponding to the symmetric ring-breathing mode of the phenyl group. To match the targeted Raman lines, the pump wavelength from the picoEmerald laser was tuned from 803.5 nm to 961 nm, covering the spectral range from the C–H stretching mode to the ring-breathing mode. The Raman pump power was set to approximately 34 mW, while the coherent Stokes beam power was maintained at around 3.75 mW to ensure a sufficient signal level. The SRS images obtained with our home-built microscopy system show good agreement with the corresponding white-light wide-field image, confirming both the spatial resolution and spectral tunability of the system.
@@ 3.1 QE-SRS on a pork muscle tissue
The Raman band at 2940 cm-1 originates from CH2/CH3 stretching modes present in both proteins and lipids and therefore serves as an indicator of overall protein–lipid content in tissue. The measurement was performed using an average optical power of 32 mW for the Raman pump and 3.75 mW for the Stokes beam focused on the sample. The white-light wide-field image is shown in Fig. 3(a). Spatial mapping of the muscle structure was achieved via raster scanning with a step size of 750 nm. The measured SRS signal (Fig. 3(b)) clearly reproduces the structural features of the muscle tissue. Fig. 3(d),(e) further illustrate the enhancement in SNR achieved by employing amplitude-squeezed light. In the case of SRS gain detection, the SNR is [7]
where is the SRS signal intensity, and are pump and Stokes beam intensities, respectively, and is the variance of the detected Stokes beam intensity that quantifies the noise level. By using amplitude-squeezed light, we are able to surpass the shot-noise level, thereby enhancing the SNR in proportion to the inverse square root of the noise reduction factor. In Fig. 3(d) and (e), we observed an SNR improvement across the entire investigated area (),
which corresponds to an approximately increase in SNR. Using this value, we can estimate the corresponding noise reduction as
which indicates a noise reduction of compared to the coherent Stokes beam configuration. The initial of squeezing was reduced to at the detector due to 24% optical loss arising from optics and sample transmission.
In addition, we investigated the scaling behavior of the SNR with respect to the Raman pump power. As discussed above, the SNR is expected to scale linearly with pump power. To validate this, we measured the SNR of the SRS signal within the white square region shown in Fig. 3(b). The results are presented in Fig. 3(c). Here, the SNR was averaged over a m2 area, while the pump power was varied from 12 mW to 49 mW with the Stokes beam power fixed at 3.75 mW. The blue and red dashed lines correspond to linear fits of the acquired data for coherent (red trace) and amplitude-squeezed (blue trace) Stokes beams, respectively. The increased slope observed in the squeezed-light case corresponds to an estimated noise reduction of approximately , yielding a 46% () improvement in SNR, which is consistent with the enhancements observed in panels (d) and (e).
@@ 3.2 Multifrequency biological imaging
To demonstrate chemical specificity, we performed SRS imaging at four vibrational bands characteristic of proteins and lipids. Explicitly, we investigated several prominent Raman vibrational modes: (general C–H stretching in proteins, lipids, and nucleic acids), widely used as a marker of total protein and lipid content; (symmetric CH2 stretching in lipids); (amide I band, primarily C=O stretching in proteins), a key indicator of protein content and secondary structure; and (CH2/CH3 bending in lipids and proteins), often used as a marker for lipid–protein balance. Together, these four bands give a complementary biochemical fingerprint of muscle tissue. For these measurements, the Raman pump was tuned to , , , and with corresponding powers at the sample of 32 mW, 32 mW, 30 mW, and 38 mW, respectively. The Stokes beam power was maintained at in all cases. However, systematically higher SNR were observed at the CH-stretch bands (2850 and 2940 cm-1) compared to the fingerprint modes. This enhancement is attributed to the larger Raman cross-sections and the higher abundance of CH groups in tissue lipids and proteins, particularly in lipid-rich structures such as cell membranes and adipose tissue. In contrast, Fig. 4 shows regions with SRS signals at 1450 cm-1 and 1650 cm-1, where the CH-stretch bands at 2850 and 2940 cm-1 are absent. This spectral pattern indicates regions with elevated protein content and minimal lipid contribution, consistent with lean myofibrillar or connective tissue in muscle [19, 2]. The results presented in Fig. 4 demonstrate a consistent enhancement in SNR across all examined Raman shifts when utilizing the amplitude-squeezed Stokes beam. A noise reduction of dB was achieved across all measured vibrational modes, corresponding to a improvement in SNR compared to the classical (coherent) Stokes configuration. Fig. 4(f) and (g) show cross-sections of the SNR maps. These cross-sections clearly illustrate the improvements achieved with squeezed light compared to coherent illumination. Along with the observed SNR enhancement, the achieved squeezing of the Stokes beam directly improves the concentration sensitivity. In particular, since the concentration sensitivity scales with the noise amplitude [7], i.e. , a noise suppression of corresponds to
which indicates that the minimum detectable concentration is reduced to of its coherent-state value, i.e., improved by approximately . This consistent enhancement underscores the robustness of quantum-assisted amplification in improving Raman signal quality. The observed gains highlight the potential of QE-SRS to extend the sensitivity of vibrational imaging in biological tissues, enabling more precise and reliable real-time biochemical analysis, even at lower optical powers or in samples with weak Raman signatures.
@@ 4 Conclusion
We validated the performance of the quantum-enhanced SRS microscope by conducting measurements on both benchmark polystyrene samples and biological samples (pork muscle tissue). The biological measurements revealed an average noise reduction of 3.6 dB, corresponding to a 51% improvement in SNR compared to classical detection. To the best of our knowledge, this represents the highest reported enhancement in QE-SRS microscopy applied to biological samples. Importantly, our results further support the applicability of squeezed-light injection in picosecond-pulsed SRS microscopy for complex biological tissue environments, while also indicating clear potential for further optimization. Along with strong squeezing, we also report, for the first time, QE-SRS across both regions in the fingerprint and high-wavenumber CH-stretch regions. Our approach provides quantum SNR enhancement in both regions, enabling sensitive and label-free imaging and identification of important biomarkers, such as proteins, lipids, amid and nucleic acids in biological samples.
Looking forward, further improvements in coupling efficiency, waveguide engineering, and pump power stabilization are expected to increase the achievable squeezing levels and long-term stability of the source.Integration with advanced lock-in detection schemes, optimized data acquisition and synchronization, and fast-scanning microscopy incorporating galvanometric scanners or resonant scanners could enable real-time, quantum-enhanced imaging of living biological samples, thereby extending the approach to practical biomedical applications. Moreover, combining squeezed-light excitation with computational imaging and machine learning approaches may open new possibilities for super-resolved Raman microscopy with reduced photo damage and enhanced molecular contrast.